# LiveScene: Language Embedding Interactive Radiance Fields for Physical Scene Rendering and Control

### v1

@@ Abstract
This paper aims to advance the progress of physical world interactive scene reconstruction by extending the interactive object reconstruction from single object level to complex scene level. To this end, we first construct one simulated and one real scene-level physical interaction dataset containing 28 scenes with multiple interactive objects per scene. Furthermore, to accurately model the interactive motions of multiple objects in complex scenes, we propose LiveScene, the first scene-level language-embedded interactive neural radiance field that efficiently reconstructs and controls multiple interactive objects in complex scenes. LiveScene introduces an efficient factorization that decomposes the interactive scene into multiple local deformable fields to separately reconstruct individual interactive objects, achieving the first accurate and independent control on multiple interactive objects in a complex scene. Moreover, we introduce an interaction-aware language embedding method that generates varying language embeddings to localize individual interactive objects under different interactive states, enabling arbitrary control of interactive objects using natural language. Finally, we evaluate LiveScene on the constructed datasets OmniSim and InterReal with various simulated and real-world complex scenes. Extensive experiment results demonstrate that the proposed approach achieves SOTA novel view synthesis and language grounding performance, surpassing existing methods by +9.89, +1.30, and +1.99 in PSNR on CoNeRF Synthetic, OmniSim #chanllenging, and InterReal #chanllenging datasets, and +65.12 of mIOU on OmniSim, respectively. Project page: https://livescenes.github.io.
Abstract
This supplementary material accompanies the main paper by providing more details for reproducibility as well as additional evaluations and qualitative results to to verify the effectiveness and robustness of LiveScene:
$\triangleright$ appendix B: Configurations of OmniSim and InterReal dataset, including scene assets, interaction variables generation, mask and prompts annotation, and dataset visualization.
$\triangleright$ appendix C: Video demonstration and anonymous link: https://livescenes. github.io.
$\triangleright$ appendix D: Additional implementation details.
$\triangleright$ appendix E: Additional experimental results, including more ablation studies, detailed view synthesis quality comparison, interactive scenes geometry comparison and language grounding comparison.
@@ 1 Introduction
Interactive objects are prevalent in our daily lives, and modeling interactable scenes from real physical world plays an essential role in various research fields, including content generation [38, 41, 40, 61, 9], animation [69, 44, 35, 42], virtual reality [59, 13, 26, 66], robotics [1, 20, 4, 21, 46], and world understanding [63, 19, 62, 18, 8]. This paper tackles the challenging and rarely explored task of reconstructing and controlling multiple interactive objects in complex scenes from a single, casually captured monocular video without previous independent modeling of geometry and kinematics.
Figure 1: Scene modeling capability, render performance, and efficiency comparisons among various methods. In contrast to static or deformable scenes, LiveScene is capable of modeling $(3+\alpha)$d interactive scenes with language grounding and achieving SOTA rendering quality on OmniSim dataset. Notably, LiveScene exhibits a significant advantage in parameter efficiency, maintaining a constant complexity regardless of interactive instance number and satisfying rendering time at $1200\times 1200$ resolution. In contrast, other methods, such as CoGS [73] and MK-Planes [54], exhibit a linear or quadratic increase in parameter quantity.
@@
Prior research on interactable scene modeling, such as CoNeRF [25] and K-Planes [54], typically adopts a joint modeling approach, combining spatial coordinates and all interaction variables as input and representing interactive scene by either implicit MLPs or feature planes. Meanwhile, CoGS [73] learns parameter offsets for different scene parts using multiple independent MLPs after establishing a 3D deformable Gaussian scene. However, these methods primarily focus on capturing interactions for a single object within a clear background, such as a single drawer, toy car, or face [76, 25, 73, 78]. As modeling extends from single objects to multiple objects in complex scenes, as shown in Figure. 1, the interaction spaces become increasingly high-dimensional and intricate, complicating these methods for accurate modeling and significantly increasing computational time and memory cost, e.g., 4$\times$ A100 GPU for 2 weeks to converge training in CoNeRF [25] and 500M Gaussian storage for a regular indoor living room in CoGS [73]. Moreover, natural language is an intuitive and necessary interface for interacting with 3D scenes, but language embedding of interactive scenes faces an even more daunting challenge: interaction variation inconsistency. For instance, methods like LERF [28], OpenNeRF [77], and GARField [29], which distill CLIP features into static 3D fields, suffer from significant failures when confronted with scene topology structure changes induced by interactions, such as the distinct structures variation of a cabinet before and after opening.
@@
To address these challenges, we propose the first scene-level language embedded interactive scenes reconstruction and control radiance fields LiveScene, which compresses the high-dimensional interaction space into compact 4D feature planes, reducing model parameters while improving optimization effectiveness. Specifically, LiveScene models the multiple object interaction space via a novel high-dimensional factorization by decomposing the interactive scene into multiple local deformable fields, which model individual interactive objects by multi-scale 4D deformable feature planes. Moreover, to achieve independent and accurate control of interactive objects, we propose a multi-scale interaction probability sampling strategy to sample the interaction-relevant 3D points in factorized local deformable fields. As for the nature language interacting interface, we introduce an interaction-aware language embedding method that generates varying language embedding to localize and control interactive objects under arbitrary interaction states, which is the first attempt to control interactive objects using nature language. Finally, we construct the first scene-level physical interaction dataset OmniSim and InterReal, containing 28 scenes with overall 70 interactive objects for evaluation. Our approach achieves state-of-the-art novel view synthesis quality, outperforming existing best methods by +9.89, +1.30, and +1.99 in PSNR on the CoNeRF Synthetic, OmniSim #chanllenging, and InterReal #chanllenging subsets, respectively. Surpassing LeRF [28], LiveScene significantly improves language embedding grounding accuracy by +65.12 of mIOU on the OmniSim dataset. Notably, our method maintains a lightweight, constant model parameter of 39M, scaling well with increasing scene complexity, as shown in Figure. 1. The contributions can be summarized as follows:
@@
  • We introduce LiveScene, the first scene-level language embedded interactive radiance field that efficiently reconstructs and controls complex physical scenes, allowing for the control of multiple interactive objects within the neural scene with diverse interaction variations and even facilitating language-based interaction.
@@
  • We propose a factorization and sampling technique that decomposes interactive scenes into local deformable fields and samples the interaction-relevant 3D points to control individual interactive objects in a complex scene. Furthermore, we introduce an interaction-relevant language embedding method that generates interaction-relevant varying language embeddings to localize and control interactive objects.
@@
  • We construct the first scene-level physical interaction dataset OmniSim and InterReal, containing 28 subsets and 70 interactive objects for evaluation. Extensive experiments demonstrate that the proposed approach achieves SOTA performance in various tasks such as novel view synthesis, video frame interpolation, and scene interactive control, showcasing robust interaction capabilities.
@@ 2 Related Work
Dynamic Scene Representation. Extending NeRF [45] to dynamic scene reconstruction has made significant progress. Related methods can generally be categorized into time-varying methods, deformable-canonical methods, and hybrid representation methods. The time-varying methods [10, 11, 37, 49, 51, 64, 75] typically model the radiance field directly over time, but struggle to separate dynamic and static objects. Deformable-canonical methods [14, 36, 50, 71] decouple dynamic deformable field and static canonical space, modeling 4D by warping points with deformable field to query the canonical features. However, these methods face challenges in scene topology changes [54]. Hybrid representation methods, on the other hand, have achieved high-quality reconstruction and fast rendering by utilizing time-space feature planes [56, 54, 5], 4D hash encoding [67], dynamic voxels [68], or triple fields [58]. Recently, several works [43, 72, 70, 32] have introduced 3D gaussians [26] into dynamic scene reconstruction, achieving high-quality real-time rendering speeds. However, these methods are limited to reconstructing dynamic scenes and lack the ability to control and understand interactive scenes.
3D Vision-language Fields. Vision-language foundational models [53, 6, 48, 24, 30] with strong generalizability and adaptability inspires numerous language embedded scene representation for 3D scene understanding [3, 80], such as open-vocabulary segmentation [79, 39, 31, 16, 65, 29], 3D visual question answering [17, 8, 23, 22], and 3D language grounding [20, 55, 7, 22]. LeRF [28] is the first to achieve open-vocabulary 3D queries by combining CLIP [53] and DINO [6] with NeRF through feature distillation. Open-NeRF [77] introduces an integrate-and-distill paradigm and leverages hierarchical embeddings to address 2D knowledge-distilling issues from SAM. LEGaussians [57] and LangSplat [52] integrate semantic features into 3D gaussians [26] and achieve precision language query and efficient rendering. However, these methods are limited to static scene understanding and fail to generalize when the interactive scene topology changes.
Controllable Scene Representation. Manipulating reconstructed assets or neural fields is of significant importance for avatar and robotic tasks [12, 2, 15, 25, 33]. CoNeRF [25] pioneered this effort by extending HyperNeRF [50] and introduce a fine-grained controlable neural field with 2D attribute mask and value annotations. CoNFies [74] proposes an automatic controllable avatar system and accelerates rendering by distilling. More recently, CoGS [73] leveraged 3D Gaussians [26] to achieve real-time control of dynamic scenes without requiring explicit control signals. However, these methods typically lack natural language interaction capabilities, relying solely on manual control. Furthermore, most works focus on single or few object interactions, disregarding the interaction between different scene parts, limiting their real-world applications.
@@ 3 Methodology
We aim to establish a representation that models $\alpha$ interactive articulated objects in a complex scene $\boldsymbol{\chi}$ from a casually captured monocular video $\{\mathbf{I}_{i}\}_{i=1}^{T}$ via rendering-based self-supervised manner. Control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ indicating object motion states, and camera poses $\{\mathbf{P}\}_{i=1}^{T}$ of each video frame are given. The overview of LiveScene is shown in Figure. 2. Section. 3.1 introduces the high-dimensional interactive space modeling and challenges. Section. 3.2 presents a multi-scale interactive space factorization and sampling strategy to compress the high-dimensional interactive space into local 4D deformable fields to model complicated interactive motions of individual objects. Section. 3.3 introduces an interaction-aware language embedding method to localize and control interactive objects using natural language.
Figure 2: The overview of LiveScene. Given a camera view and control variable $\boldsymbol{\kappa}$ of one specific interactive object, a series 3D points are sampled in a local deformable field that models the interactive motions of this specific interactive object, and then the interactive object with novel interactive motion state is generated via volume-rendering. Moreover, an interaction-aware language embedding is utilized to localize and control individual interactive objects using natural language.
@@ 3.1 Interactive Space
Assuming a non-rigidly interactive scene $\boldsymbol{\chi}$ with $\alpha$ control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ corresponding to $\alpha$ objects, we delineate its representation by a high-dimensional function:
@@
where $\rho$ is the model of the representation, $\mathbf{x}\in\mathbb{R}^{3}$ are spatial coordinates, $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ are control variables, $\mathcal{H}$ is a set of optional additional parameters (e.g., the view direction), and $\boldsymbol{\theta}$ stores the scene information. The function outputs scene properties $\mathbf{y}$ for the given position $\mathbf{x}$ and $\boldsymbol{\kappa}$ sampling from a ray $\mathbf{r}$, where $\mathbf{y}$ can be represented with color, occupancy, signed distance, density, and BRDF parameters. In this paper, we focus on color, probability, and language embedding. Distinguishing from 3d static scene or 4d dynamic scene modeling, the sampling point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$ in interactive scene is high-dimensional and variable in topological structure, complicating the scene feature storage $\boldsymbol{\theta}$ and the optimization of representation model $\rho$, leading to significant time-consuming or memory-intensive training in [25, 73].
@@ 3.2 Multi-scale Interaction Space Factorization
For a $(3+\alpha)$-dimensional interactive space $\boldsymbol{\chi}=\mathbb{R}^{(3+\alpha)}$ containing $\alpha$ control variables, we aim to explicitly represent the high-dimensional space in a concise and compact storage, thereby reduce memory usage and improve optimization. As illustrated in Figure. 2, interactive objects exhibit mutual independence, and interaction features $\mathcal{F}_{\alpha}\in\boldsymbol{\theta}$ unevenly distribute in the $(3+\alpha)$-dimensional interactive space $\boldsymbol{\chi}$ and aggregate into cluster centers $\mathbf{c}_{i},i\leq\alpha$. Thus, there exists a set of hyperplanes that partition the space into disjoint regions =$\mathcal{R}=\left\{\mathcal{R}_{1},\mathcal{R}_{2},\cdots\mathcal{R}_{\alpha}\right\}$, with each subregion $\mathcal{R}_{i}$ containing a local 4D deformable field $\boldsymbol{\theta}_{d}$, as shown in Figure. 3. Hence, the interaction features $\boldsymbol{\theta}(\mathbf{p})$ at sampling point $\mathbf{p}\in\mathbb{R}^{3+\alpha}$ can be projected into a compact 4-dimensional space $\mathbb{R}^{4}$ in Figure. 3 by a transformation $\mathbf{T}$:
$\Theta(\boldsymbol{\theta}(\mathbf{p}),\mathbf{T})\to\boldsymbol{\theta}_{d}(\mathbf{p_{d}}),\quad\mathbf{p}\in\mathbb{R}^{(\alpha+3)},\quad\mathbf{p_{d}}\in\mathbb{R}^{4},$ (2)
where $\Theta$ represents the operation of projecting high-dimensional features into a 4-dimensional space, and $\mathbf{T}$ denotes the corresponding projection matrix.
Multi-scale Interactive Ray Sampling. Solving $\Theta;\mathbf{T}$ is complex and unwise, we consider using ray sampling to perform the projection. As shown in Figure. 2, assuming a ray $\mathbf{r}(t)=\left[\mathbf{o_{x}}|\mathbf{o}_{\boldsymbol{\kappa}}\right]+t\left[\mathbf{d_{x}}|\mathbf{d}_{\boldsymbol{\kappa}}\right]$ with origin $\mathbf{o_{x}}\in\mathbb{R}^{3},\mathbf{o}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, and direction $\mathbf{d_{x}}\in\mathbb{R}^{3},\mathbf{d}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, the ray intersects with the interaction instances =$\mathcal{R}=\left\{\mathcal{R}_{1},\mathcal{R}_{2},\cdots\mathcal{R}_{\alpha}\right\}$ at a point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$, where $\mathbf{x}\in\mathbb{R}^{3}$ is the 3D position and $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ is the interaction variables. For a given intersection point $\mathbf{p}$, the deformable features $\mathcal{F}_{d}$ can be retrieved from the corresponding local 4D deformable field $\boldsymbol{\theta}_{d}$ by:
$\mathcal{F}_{d}=\boldsymbol{\theta}_{d}(\mathbf{p_{d}}),\quad\mathbf{p_{d}}=\left[\mathbf{x}|\boldsymbol{\kappa}_{d}\right],\quad d=\operatorname*{arg\,max}_{i}\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}_{s}(\mathbf{x}))_{i},$ (3)
where $\boldsymbol{\theta}_{s}(\mathbf{x})$ is the probability features at position $\mathbf{x}$ from 3d feature planes, and $\operatorname*{arg\,max}_{i}\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}_{s}(\mathbf{x}))_{i}$ maps the interaction variables $\boldsymbol{\kappa}$ to the most probable cluster region $\mathcal{R}_{d}$ in the 4D space, and can be optimized by minimizing the focal loss of mask $\tilde{\mathbf{y}}_{m}$ across all the training camera views and images: $\underset{\Theta,\boldsymbol{\theta}_{s}}{\arg\min}~{}\mathcal{L}_{\text{focal}}(\tilde{\mathbf{y}}_{m},\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}_{s}(\mathbf{x})))$ in supplemental appendix D. Next, the deformable features $\mathcal{F}_{d}$ are used to compute the local deformable scene properties $\mathbf{y}$ at the sampling point $\mathbf{p}$. In this way, the high-dimensional interaction features $\mathcal{F}_{\boldsymbol{\alpha}}$ are factorized into a 4D space $\mathcal{F}_{d}$ by a lightweight transformation $\Theta$ modeling with interaction probability decoder and 3d feature planes in Figure. 2, supervised by deformable masks $\tilde{\mathbf{y}}_{m}$. Moreover, leveraging K-Planes [54], the multiple 4d local deformable space can be further compressed in only $\mathbf{C}_{4}^{2}=6$ feature planes. We iteratively sample from coarse to fine within the multi-scale feature plane, retrieving the maximum probability interaction variables $\boldsymbol{\kappa}_{d}$ and indices $d$ at each scale.
@@
Figure 3: Illustration of the proposed hyperplanar factorization for compact storage and efficient rendering. We maintain mutiple local deformable fields $\mathcal{F}_{d}$ for each interactive object =$\mathcal{R}=\left\{\mathcal{R}_{1},\mathcal{R}_{2},\cdots\mathcal{R}_{\alpha}\right\}$ in the 4D space, and project the high-dimensional interaction features $\mathcal{F}_{\alpha}$ into a compact 4D space $\mathcal{F}_{d}$, which can be further compressed into multiscale feature planes.
@@
Feature Repulsion and Probability Rejection. A latent issue is that optimizing the interaction probability decoder with varying masks can lead to blurred boundaries in the local deformable field, further causing ray sampling and feature storage conflicts. As illustrated in Figure. 4(a), consider two adjacent local deformable regions $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$, and a point $\mathbf{p}$ in high-dimensional space, suppose $\mathbf{p}$ moves from the cluster center $\mathbf{c}_{i}$ of $\mathcal{R}_{i}$ towards the cluster center $\mathbf{c}_{j}$ of $\mathcal{R}_{j}$, then the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{i}$ gradually decreases, while the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{j}$ increases. To avoid sampling conflicts and feature oscillations at the boundaries, we introduce a repulsion loss to amplify the feature differences between distinct deformable scenes, thereby promoting the separation of deformable field:

### v2

@@ Abstract
This paper scales object-level reconstruction to complex scenes, advancing interactive scene reconstruction. We introduce two datasets, OmniSim and InterReal, featuring 28 scenes with multiple interactive objects. To tackle the challenge of inaccurate interactive motion recovery in complex scenes, we propose LiveScene, a scene-level language-embedded interactive radiance field that efficiently reconstructs and controls multiple objects. By decomposing the interactive scene into local deformable fields, LiveScene enables separate reconstruction of individual object motions, reducing memory consumption. Additionally, our interaction-aware language embedding localizes individual interactive objects, allowing for arbitrary control using natural language. Our approach demonstrates significant superiority in novel view synthesis, interactive scene control, and language grounding performance through extensive experiments. Project page: https://livescenes.github.io.
Figure 1: LiveScene enables scene-level reconstruction and control with language grounding. Left: Language-interactive articulated object control in Nerfstudio. Right: LiveScene achieves SOTA rendering quality on OmniSim dataset and exhibits a significant advantage in parameter efficiency.
@@ 1 Introduction
Interactive objects are prevalent in our daily lives, and modeling interactable scenes from the real physical world plays an essential role in various research fields, including content generation [33, 35, 52], animation [58, 37, 30], virtual reality [50, 9, 32], robotics [1, 61, 44], and world understanding [16, 14, 53]. This paper tackles the challenging and rarely explored task of reconstructing and controlling multiple interactive objects in complex scenes from a single, casually captured monocular video without previous independent modeling of geometry and kinematics. Prior research on interactable scene modeling, such as CoNeRF [20] and K-Planes [10], typically adopts a joint modeling approach, combining spatial coordinates and all interaction variables as input and representing interactive scene by either implicit MLPs or feature planes. Meanwhile, CoGS [63] learns parameter offsets for different scene parts using multiple independent MLPs after establishing a 3D deformable Gaussian scene. However, these methods primarily focus on capturing interactions for a single object within a clear background, such as a single drawer, toy car, or face [66, 20, 63, 68]. As modeling extends from single objects to multiple objects in complex scenes, as shown in Figure. 1, the interaction spaces become increasingly high-dimensional, complicating these methods for accurate modeling and significantly increasing computational time and memory cost, e.g., 4$\times$ A100 GPU for 2 weeks to converge training in CoNeRF [20] and 500M Gaussian storage for a regular indoor living room in CoGS [63]. Moreover, natural language is an intuitive and necessary interface for interacting with 3D scenes, but language embedding of interactive scenes faces an even more daunting challenge: interaction variation inconsistency. For instance, methods like LERF [23], and OpenNeRF [67], which distill CLIP features into static 3D fields, suffer from significant failures when confronted with scene topology structure changes induced by interactions, such as the distinct structures variation of a cabinet before and after opening.
@@
To address these challenges, we propose LiveScene, the first scene-level language-embedded radiance fields, which compresses high-dimensional interaction spaces into compact 4D feature planes, reducing model parameters while improving optimization effectiveness. LiveScene models multiple object interactions via novel high-dimensional factorization, decomposing the scene into local deformable fields that model individual objects with multi-scale 4D deformable feature planes. To achieve independent control, we introduce a multi-scale interaction probability sampling strategy for factorized local deformable fields. Our interaction-aware language embedding method generates varying language embeddings to localize and control objects under arbitrary states, enabling natural language control. Finally, we construct the first scene-level physical interaction datasets, OmniSim and InterReal, featuring 28 scenes with 70 interactive objects for evaluation.
Experiment results show that our approach achieves SOTA novel view synthesis quality, outperforming existing best methods by +9.89, +1.30, and +1.99 in PSNR on the CoNeRF Synthetic, OmniSim #chanllenging, and InterReal #chanllenging subsets, respectively. Surpassing LeRF [23], LiveScene significantly improves language grounding accuracy by +65.12 of mIOU on the OmniSim dataset. Notably, our method maintains a lightweight, constant model parameter of 39M, scaling well with increasing scene complexity, as shown in Figure. 1. Contributions can be summarized as:
@@
  • We propose LiveScene, the first scene-level language-embedded interactive radiance field, which efficiently reconstructs and controls complex physical scenes, enabling manipulation of multiple articulated objects and language-based interaction.
@@
  • We propose a factorization technique that decomposes interactive scenes into local deformable fields and samples relevant 3D points, enabling control of individual objects. Additionally, we introduce an interaction-aware language embedding method that generates varying embeddings, allowing for language-based control and localization.
@@
  • We construct the first scene-level physical interaction dataset OmniSim and InterReal, containing 28 subsets and 70 interactive objects for evaluation. Extensive experiments demonstrate our SOTA performance and robust interaction capabilities.
@@ 2 Related Work
Dynamic Scene Representation. Extending NeRF [38] to dynamic scene reconstruction has made significant progress. Related methods can generally be categorized into time-varying methods, deformable-canonical methods, and hybrid representation methods. The time-varying methods [7, 41, 54, 65] typically model the radiance field directly over time, but struggle to separate dynamic and static objects. Deformable-canonical methods [11, 31, 42, 60] decouple dynamic deformable field and static canonical space, modeling 4D by warping points with deformable field to query the canonical features. However, these methods face challenges in scene topology changes [10]. Hybrid representation methods, on the other hand, have achieved high-quality reconstruction and fast rendering by utilizing time-space feature planes [47, 10, 3], 4D hash encoding [56], dynamic voxels [57], or triple fields [49]. Recently, several works [36, 62, 59, 27] have introduced 3D gaussians [21] into dynamic scene reconstruction, achieving high-quality real-time rendering speeds. However, these methods are limited to reconstructing dynamic scenes and lack the ability to control and understand interactive scenes.
3D Vision-language Fields. Vision-language foundational models [45, 40, 25] with strong generalizability and adaptability inspires numerous language embedded scene representation for 3D scene understanding [2, 70], such as open-vocabulary segmentation [34, 13, 55, 24], 3D visual question answering [15, 6, 19, 18], and 3D language grounding [17, 46, 5, 18]. LeRF [23] is the first to achieve open-vocabulary 3D queries by combining CLIP [45] and DINO [4] with NeRF through feature distillation. Open-NeRF [67] introduces an integrate-and-distill paradigm and leverages hierarchical embeddings to address 2D knowledge-distilling issues from SAM. LEGaussians [48] and LangSplat [43] integrate semantic features into 3D gaussians [21] and achieve precision language query and efficient rendering. However, these methods are limited to static scene understanding and fail to generalize when the interactive scene topology changes.
Controllable Scene Representation. Manipulating reconstructed assets or neural fields is of significant importance for avatar and robotic tasks [8, 12, 20, 28]. CoNeRF [20] pioneered this effort by extending HyperNeRF [42] and introduce a fine-grained controlable neural field with 2D attribute mask and value annotations. CoNFies [64] proposes an automatic controllable avatar system and accelerates rendering by distilling. More recently, CoGS [63] leveraged 3D Gaussians [21] to achieve real-time control of dynamic scenes without requiring explicit control signals. However, these methods typically lack natural language interaction capabilities, relying solely on manual control. Furthermore, most works focus on single or few object interactions, disregarding the interaction between different scene parts, limiting their real-world applications.
@@ 3 Methodology
Figure 2: The overview of LiveScene. Given a camera view and control variable $\boldsymbol{\kappa}$ of one specific interactive object, a series of 3D points are sampled in a local deformable field that models the interactive motions of this specific interactive object, and then the interactive object with novel interactive motion state is generated via volume-rendering. Moreover, an interaction-aware language embedding is utilized to localize and control individual interactive objects using natural language.
We aim to establish a representation that models $\alpha$ interactive articulated objects in a complex scene from a monocular video via a rendering-based self-supervised manner. Control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ indicating object motion states and camera poses of each video frame are given. The overview of LiveScene is shown in Figure. 2. Section. 3.1 introduces the high-dimensional interactive space modeling and challenges. Section. 3.2 presents a multi-scale interactive space factorization and sampling strategy to compress the high-dimensional interactive space into local 4D deformable fields, and model complicated interactive motions of individual objects. Section. 3.3 introduces an interaction-aware language embedding method to localize and control interactive objects using natural language.
@@ 3.1 Interactive Space
Assuming a non-rigidly interactive scene with $\alpha$ control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ corresponding to $\alpha$ objects, we delineate its representation by a high-dimensional function:
@@
where $\rho$ is the model of the representation, $\mathbf{x}\in\mathbb{R}^{3}$ are spatial coordinates, $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ are control variables, $\mathcal{H}$ is a set of optional additional parameters (e.g., the view direction), and $\boldsymbol{\theta}$ stores the scene information. The function outputs scene properties $\mathbf{y}$ for the given position $\mathbf{x}$ and $\boldsymbol{\kappa}$ sampling from a ray $\mathbf{r}$, where $\mathbf{y}$ can be represented with color, occupancy, signed distance, density, and BRDF parameters. This paper focuses on color, probability, and language embedding. Distinguishing from 3d static scene or 4d dynamic scene modeling, the sampling point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$ in the interactive scene is high-dimensional and variable in topological structure, complicating the scene feature storage and the optimization of representation model, leading to significant time-consuming or memory-intensive training in [20, 63].
@@ 3.2 Multi-scale Interaction Space Factorization
For the $(3+\alpha)$-dimensional interactive space containing $\alpha$ control variables, we aim to explicitly represent the high-dimensional space in a concise and compact storage, thereby reducing memory usage and improving optimization. As illustrated in Figure. 2 and Figure. 10, objects exhibit mutual independence, and interaction features are distributed in the $(3+\alpha)$-dimensional interactive space and aggregate into cluster centers. Thus, there exists a set of hyperplanes that partition the space into disjoint regions, with each region containing a local 4D deformable field, as shown in Figure. 3. Hence, the interaction features at sampling point $\mathbf{p}\in\mathbb{R}^{3+\alpha}$ can be projected into a compact 4-dimensional space $\mathbb{R}^{4}$ in Figure. 3 by a transformation.
@@
Figure 3: Illustration of hyperplanar factorization for compact storage. We maintain multiple local deformable fields for each interactive object region $\mathcal{R}_{i}$, and project high-dimensional interaction features into a compact 4D space, which can be further compressed into multiscale feature planes.
Multi-scale Interactive Ray Sampling. We consider using ray sampling to perform the projection transformation. As shown in Figure. 3, assuming a ray $\mathbf{r}(t)=\left[\mathbf{o_{x}}|\mathbf{o}_{\boldsymbol{\kappa}}\right]+t\left[\mathbf{d_{x}}|\mathbf{d}_{\boldsymbol{\kappa}}\right]$ with origin $\mathbf{o_{x}}\in\mathbb{R}^{3},\mathbf{o}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, and direction $\mathbf{d_{x}}\in\mathbb{R}^{3},\mathbf{d}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, the ray intersects with the interaction region at a point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$, where $\mathbf{x}\in\mathbb{R}^{3}$ is the 3D position and $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ is the interaction variables. For a given intersection point $\mathbf{p}$, the deformable features can be retrieved from the corresponding local 4D deformable field by maximizing sampling probability $\mathbf{P}$:
$\mathbf{p}_{u}=\left[\mathbf{x}|\boldsymbol{\kappa}(u)\right],\quad u=\operatorname*{arg\,max}_{i}\{\mathbf{P}_{i}\},\quad\mathbf{P}=\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}),$ (2)
where $\mathbf{p}_{u}$ and $\boldsymbol{\theta}$ are the 4D sampling point and probability features at position $\mathbf{x}$ from 3D feature planes. The maximum argument operation of probability decoder $\Theta(\boldsymbol{\kappa},\boldsymbol{\theta})$ maps the interaction variables $\boldsymbol{\kappa}$ to the most probable cluster region in the 4D space, and can be optimized by minimizing the focal loss $\mathcal{L}_{\text{focus}}$ of mask across all the training camera views:
$\mathcal{L}_{\text{focus}}=\beta\cdot\left(1-e^{\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})}\right)^{\gamma}\cdot\left(-\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})\right),$ (3)
where $\mathbf{M}$ is the ground truth mask label, $\hat{\mathbf{P}}$ is the probability map rendering from the interactive probability field, $\beta$ is the balancing factor, and $\gamma$ is the focusing parameter.
Next, the deformable features are used to render local deformable scene color at the sampling point $\mathbf{p}$. In this way, the high-dimensional interaction features are factorized into a 4D space by a lightweight transformation modeling with interaction probability decoder and 3D feature planes in Figure. 2, supervised by deformable masks $\mathbf{M}$. Moreover, leveraging K-Planes [10], the multiple 4D local deformable space can be further compressed in only $\mathbf{C}_{4}^{2}=6$ feature planes. We iteratively sample from coarse to fine within the multi-scale feature plane, retrieving the maximum probability interaction variables $\boldsymbol{\kappa}_{u}$ and indices $u$ at each scale.
@@
Feature Repulsion and Probability Rejection. A latent challenge is that optimizing the interaction probability decoder with varying masks can lead to blurred boundaries in the local deformable field, further causing ray sampling and feature storage conflicts. As illustrated in Figure. 4(a), consider two adjacent local deformable regions $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$, and a point $\mathbf{p}$ in high-dimensional space, suppose $\mathbf{p}$ moves from the cluster center of $\mathcal{R}_{i}$ towards the cluster center of $\mathcal{R}_{j}$, then the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{i}$ gradually decreases, while the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{j}$ increases. To avoid sampling conflicts and feature oscillations at the boundaries, we introduce a repulsion loss for ray pairs $(\mathbf{r}_{i},\mathbf{r}_{j})$, and amplify the feature differences between distinct deformable regions, promoting the separation of deformable field: